OpenStudy (anonymous):

Please help :) The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08. A. 1.5 B. 2 C. 2.4 D. 3

Directrix (directrix):

A = P(1 + rt) 168 = 150 ( 1 + .08)*t @alex8699 Can you write the next line?

OpenStudy (anonymous):

i knew how to write the equation out, can't figure out what t is

OpenStudy (anonymous):

168 = 150*( 1 + t * 0,08 )?

OpenStudy (anonymous):

1.5?

Directrix (directrix):

Correct.

OpenStudy (anonymous):

yeah just solved it, thanks though :) @FINYA

Directrix (directrix):

You can use the simple interest calculator at http://www.calculatorsoup.com/calculators/financial/simple-interest-plus-principal-calculator.php

OpenStudy (anonymous):

t = (168/150 - 1)/0.08

Directrix (directrix):

>>1.5? Yes. @alex8699